OpenStudy (anonymous):

How to simplify this? \[ \frac{-\sqrt{3} - 1}{1+(-\sqrt{3})*1} \]

OpenStudy (mathmate):

I thought @TheSmartOne solved it earlier !?

OpenStudy (mathmate):

Hint: use identities \(a^2-b^2 = (a+b)(a-b), and\) \((a+b)^2 = a^2+2ab+b^2\) then rationalize the denominator.

OpenStudy (anonymous):

\[ \frac{-\sqrt{3} - 1}{1+(-\sqrt{3})*1}\] \[ \frac{-\sqrt{3} - 1}{1-\sqrt{3}}\] \[ \frac{-\sqrt{3} - 1}{1-\sqrt{3}} * \frac{1-\sqrt{3}} {1-\sqrt{3}} \] \[ \frac{2}{2(2-\sqrt{3})} \] Is this correct??